An experienced worker can paint a room three times faster than her apprentice can. Working together, they can do the job in 6 hours. How long would it take the experienced worker to finish the job if she worked alone?
Answer by jorel1380(3719)

1/x+1/3x=1/6
6+2=x
8=x
It takes the experienced worker 8 hours working alone----the apprentice takes 24 hours alone.
1/8+1/24=3/24+1/24=4/24=1/6..